# An Equation of State for Turbulence in the Gross-Pitaevskii model

### v1

@@ Abstract
Having verified that the energy input (at low $k$) and dissipation (at high $k$) rates are equal, we turn to the direct calculation of the scale-resolved energy flux $\Pi_{\varepsilon}(k)$ and particle flux $\Pi_{N}(k)$ (see [27] for the formal definition of the fluxes). As shown in Fig. 2(c), $\Pi_{\varepsilon}$ is momentum independent in steady state and equal to the energy injection rate $\epsilon_{\textrm{in}}$. The energy injected at low $k$ is transported to higher $k$ by the momentum-independent flux $\Pi_{\varepsilon}(k)$ and dissipated at $\gtrsim k_{\textrm{D}}$. As expected for an energy cascade, $\Pi_{\varepsilon}$ is also (nearly) independent of the dissipation scale $k_{\textrm{D}}$  [^8] Note that the slow decrease of the plateau of $\Pi_{\varepsilon}$ is due to a progressive depletion of the condensate, which results in a (slow) decrease of $\epsilon_{\textrm{in}}$ [corresponding horizontal dashed lines in the upper panel of Fig. 2(c)].. On the other hand, $\Pi_{N}$ is also momentum independent in steady state, but its plateau value decreases as $\propto k_{\textrm{D}}^{-2}$, as expected for particles with a quadratic dispersion relation [see dotted line in the lower panel of Fig. 2(c)]  [^9] Note that it is only in the limit $k_{\textrm{D}}\rightarrow\infty$ that one recovers the expected limits $\Pi_{\varepsilon}\rightarrow\epsilon$ with $\Pi_{N}\rightarrow 0$ for the direct energy cascade [20].. We define $\epsilon$, the scale-invariant energy-density flux, as the plateau value of $\Pi_{\varepsilon}$; in the rest of the paper we compute $\epsilon$ as $\langle Fvn\rangle$.
@@
Finally, we investigate the relation between the two far-from-equilibrium state variables of the system, $\epsilon$ and $n_{0}$. Fig. 3(a) shows $n_{0}$ and $\epsilon$ for different $a$; we express both state variables using the instantaneous natural scales of the GP equation $\xi_{t}$, $\tau_{t}$, and $\zeta_{t}$ defined through the instantaneous density $n$  [^10] Note that $\xi_{t}$ is not directly present in the axes of Fig. 3, but is used in the calculation of $n_{0}=N_{k}k^{-3}(k\xi_{t})^{-0.5}$.. Data for different values of $a$ fall on a (nearly) universal curve, demonstrating that the only relevant scales describing the turbulent cascade are the intrinsic scales of the universal GP model Eq. (1) and all the dependence on the system and drive scales ($L$, $R$, $U_{\textrm{s}}$, $\omega$, $k_{\textrm{D}}$) drops out  [^11] The independence of our EoS with $k_{\textrm{D}}$ cannot rigorously hold as for large enough $k_{\textrm{D}}$ (presumably beyond what we explored), WWT results should be recovered. .
@@
We find that the GP model contains a turbulent EoS that is a power law of the form $n_{0}/n=C(\tau_{t}\epsilon/[n\zeta_{t}])^{b}$, with $C=29(2)$ and $b=0.67(2)$, which is not described by known paradigms of turbulence. In particular, it cannot be explained by $4$-wave processes: the relation between $n_{0}$ and $\epsilon$ would have to be independent of the total density $n$ which is not the case for the EoS constructed here  [^12] The factor of $\xi^{0.5}$ on the $y$-axis is a finite-$k_{\textrm{D}}$ effect, but even in its absence our EoS would depend on $n$; within the GP model, only the WWT prediction of a power-law EoS with an exponent of $1/3$ is independent of $n$.. Furthermore, the power-law exponent $0.67(2)$ is inconsistent with any wave-kinetic description, as a kinetic theory with $\ell$-wave interactions predicts an exponent $b=1/(\ell-1)\leq 0.5$. Interestingly, this exponent for the amplitude of mode occupation is numerically the same as the exponent for the amplitude of the energy spectrum for hydrodynamic turbulence $E_{\text{K41}}\propto\epsilon^{2/3}$ [42]. However, this is likely coincidental. First, these quantities are different; one step further: the energy spectrum associated with our $N_{k}$ is $E(k)\propto k^{0.5}$ for $\gamma=3.5$ (assuming a quadratic dispersion relation), while $E_{\text{K41}}\propto k^{-5/3}$. Secondly, the predictions of K41 are independent of interactions, while our universal EoS strongly depends on them (as the rescaling of $\epsilon$ with $\zeta_{t}/\tau_{t}$ corresponds to a rescaling with $a^{2}$).
@@
Finally, we note that the slow decrease of $N$ due to evaporative losses above $k_{\textrm{D}}$ implements a slow thermodynamic-like process. Indeed, as $N$ decreases, the state variable $\epsilon\propto N$ also decreases, resulting in a slow change of the far-from-equilibrium state (see also [27]). In Fig. 3(b), we show that the simultaneous change of the state variables $\epsilon$ and $n_{0}$ follows the EoS: when rescaled to the instantaneous GP scales, the instantaneous $\epsilon$ and $n_{0}$ ‘slide up’ on our universal EoS line over time, see blue to red shades  [^13] As shown in [27], $n_{0}$ and $\epsilon/n$ are roughly constant over time, resulting in the increase of $n_{0}/n$ and $\tau_{t}\epsilon/(n\zeta_{t})$. This is reminiscent of the concept of thermodynamic quasi-static processes, for which infinitesimal changes of external constraints take a system through a dense succession of equilibrium states [44]. Remarkably, we find that this concept generalizes to states that are even locally far from equilibrium.
@@
In conclusion, we numerically investigated the properties of a turbulent cascade arising in the GP model when a box trapped gas is periodically driven and showed that it can be described by a universal EoS relating the turbulent state variables, the energy flux and the cascade amplitude. The form of our EoS is inconsistent with any kinetic theory, posing a new theoretical challenge. The comparision of our mean-field simulations to experimental measurements provides valuable benchmarks for testing the validity of mean-field theories in far-from-equilibrium scenarios. Note that even though our work provides some elements to understand the experimental results of [13], the experimental EoS cannot be fully understood solely within the framework of the GP model as the data features a scaling with $na^{3}$ that is incompatible with the universal GP model (see also [27]). In the future, it would be interesting to study on its own the incompressible parts of this turbulent scenario [19]. One should also investigate the fluctuations of $N_{k}$; comparing these fluctuations between experiments and classical field simulations could shed light on the role of quantum fluctuations in turbulent quantum fluids.

### v2

@@ Abstract
Having verified that the energy input rate (at low $k$) and dissipation rate (at high $k$) are equal, we turn to the direct calculation of the scale-resolved energy flux $\Pi_{\varepsilon}(k)$ and particle flux $\Pi_{N}(k)$ (see [36] for the formal definition of the fluxes). As shown in Fig. 2(c) in the steady state, $\Pi_{\varepsilon}$ is momentum independent in the inertial range, a certain range of momenta where neither forcing nor dissipation takes place; this flux transports the energy injected at low $k$ to higher $k$, up to $k_{\textrm{D}}$. Importantly, $\Pi_{\varepsilon}$ in the inertial range is equal to $\epsilon_{\textrm{in}}$. As expected for an energy cascade, $\Pi_{\varepsilon}$ is also (nearly) independent of the dissipation scale $k_{\textrm{D}}$  [^9] Note that the slow decrease of the plateau of $\Pi_{\varepsilon}$ is due to a progressive depletion of the condensate, which results in a (slow) decrease of $\epsilon_{\textrm{in}}$ [corresponding horizontal dashed lines in the upper panel of Fig. 2(c)].. On the other hand, $\Pi_{N}$ is also momentum independent in steady state, but its plateau value decreases as $\propto k_{\textrm{D}}^{-2}$, as expected for particles with a quadratic dispersion relation [see dotted line in the lower panel of Fig. 2(c)]  [^10] Note that it is only in the limit $k_{\textrm{D}}\rightarrow\infty$ that one recovers the expected limits $\Pi_{\varepsilon}\rightarrow\epsilon$ with $\Pi_{N}\rightarrow 0$ for the direct energy cascade [56].. We define $\epsilon$, the scale-invariant energy-density flux, as the plateau value of $\Pi_{\varepsilon}$; in the rest of the paper we compute $\epsilon$ as $\langle Fvn\rangle$.
Figure 3: A universal equation of state of the GP model. (a) The numerical calculations for different $a$ collapse onto a universal curve when the state variables $\epsilon$ and $n_{0}$ are expressed in the instantaneous natural scales of the GP model ($n$, $\xi_{t}$, $\zeta_{t}$, $\tau_{t}$). The dashed line is a power-law fit to the data that gives $n_{0}/n=29(2)(\tau_{t}\epsilon/[n\zeta_{t}])^{0.67(2)}$, and the pink band shows the fit uncertainty. (b) The comparison of our numerical EOS (gray symbols) with the experimental data from [14] (colored symbols). The color coding of the experimental data is based on the gas parameter $na^{3}$; note that the experimental fluxes are multiplied by $1.3$ compared to the data of [14] to account for $\alpha\neq 1$. The vertical error bars of the experimental data represent the uncertainties due to different $\gamma_{0}$ in simulations and experiments (see [36]).
@@
EOS for turbulence. Finally, we investigate the relation between the two far-from-equilibrium state variables of the system, $\epsilon$ and $n_{0}$. Fig. 3(a) shows $n_{0}$ and $\epsilon$ for different $a$; we express both state variables using the instantaneous natural scales of the GP equation $\xi_{t}$, $\tau_{t}$, and $\zeta_{t}$ defined through the instantaneous density $n$  [^11] Note that $\xi_{t}$ is not directly present in the axes of Fig. 3, but is used in the calculation of $n_{0}=N_{k}k^{-3}(k\xi_{t})^{-0.5}$.. Data for different values of $a$ fall on a (nearly) universal curve, demonstrating that the only relevant scales describing the turbulent cascade are the intrinsic scales of the universal GP model Eq. (1) and all the dependence on the system and drive scales ($L$, $R$, $U_{\textrm{s}}$, $\omega$, $k_{\textrm{D}}$) drops out  [^12] The independence of our EOS with $k_{\textrm{D}}$ presumably will not hold for large enough $k_{\textrm{D}}$ (beyond what we studied), as it is reasonable to expect that as $k\rightarrow\infty$ (with $k<k_{\textrm{D}}$), WWT results should be recovered..
@@
The data of Fig. 3(a) show that the GP model contains a turbulent EOS that is a power law of the form $n_{0}/n=C(\tau_{t}\epsilon/[n\zeta_{t}])^{b}$, with $C=29(2)$ and $b=0.67(2)$, which is not described by known paradigms of turbulence. The two broad paradigms are wave (compressible-energy dominated) and vortex (incompressible-energy dominated) turbulence. Our $b$ is inconsistent with any wave-kinetic description, as a kinetic theory with $\ell$-wave interactions predicts an exponent $b=1/(\ell-1)\leq 0.5$. Furthermore, the most plausible $4$-wave theory predicts that the relation between $n_{0}$ and $\epsilon$ would have to be independent of the total density $n$, which is not the case for the EOS constructed here  [^13] For an extensive comparison of our EOS to the prediction of $4$-wave theory (as well as how the respective $N_{k}$ can be deduced), see Section VII and Fig. S3141414The factor of $\xi^{0.5}$ on the $y$-axis is a finite-$k_{\textrm{D}}$ effect, but even in its absence our EOS would depend on $n$; within the GP model, only the WWT prediction of a power-law EOS with an exponent of $1/3$ is independent of $n$..
To go further, we decompose in [36] the energy spectrum into compressible and incompressible parts [25] and show that while the compressible part is typically larger than the incompressible one in the $k$ range used to extract $n_{0}$, the two are of the same magnitude; moreover, the incompressible part increases more rapidly with $\epsilon$ (similar observations were also recently reported in [21]). The observed EOS is therefore likely a result of an interplay between compressible (wave) and incompressible (vortex) excitations. As far as we are aware, there is no existing theory that describes this regime of ‘mixed’ quantum turbulence and the shape of the EOS remains to be understood. We further note that the value of $b$ being the same as the exponent of the scaling of the energy spectrum for vortex (i.e. hydrodynamic) turbulence $E_{\text{K41}}\propto\epsilon^{2/3}$ [51] is likely coincidental, as this prediction is for the incompressible energy spectrum while our EOS is for the momentum distribution; besides, our computed incompressible energy spectra are also distinct from the K4141 prediction [36]. We note, however, that Kolmogorov turbulence can also be realized and measured in ultracold-atom experiments [17] (for $k\xi\ll 1$) but in that case the K4141 spectrum originates from (coarse-grained) velocity field rather than the (atomic) momentum distribution.
Figure 4: Quasi-static process far from equilibrium. The instantaneous far-from-equilibrium state variables for different $t$ follow the EOS, in analogy with a quasi-static process in equilibrium thermodynamics. The data correspond to $a=100a_{0}$ and the blue star here corresponds to the green star in Fig. 3(a). The inset shows the fraction of particles $N/N_{0}$ left in the system.
In Fig. 3(b) we compare our calculations (grey symbols) with the experimental measurements [14] (colored symbols); see also [36] for details. The experimental data are above the numerical results. Note that the experimental data do not represent a universal EOS on this plot: they collapse onto a single curve when both axes are rescaled with empirically determined powers of the gas parameter $na^{3}$ [14], a scaling that is inherently beyond the GP model. However, we note that in the limit $n\rightarrow\infty$ and $na^{3}\rightarrow 0$ the GP model is expected to be a good description of the experimental setting and Fig. 3(b) shows that the experimental data appear to approach our numerical results for lower $na^{3}$.
@@
Finally, we note that the slow decrease of $N$ due to evaporative losses above $k_{\textrm{D}}$ implements a slow thermodynamic-like process. Indeed, as $N$ decreases, the state variable $\epsilon\propto N$ also decreases, resulting in a slow change of the far-from-equilibrium state (see also [36]). In Fig. 4, we show that the simultaneous change of the state variables $\epsilon$ and $n_{0}$ follows the EOS: when rescaled to the instantaneous GP scales, the instantaneous $\epsilon$ and $n_{0}$ ‘slide up’ on our universal EOS line over time, see blue to red shades  [^15] As shown in [36], $n_{0}$ and $\epsilon/n$ are roughly constant over time, resulting in the increase of $n_{0}/n$ and $\tau_{t}\epsilon/(n\zeta_{t})$. The points in Fig. 3(a) correspond to shaking times $\approx 2t_{\textrm{D}}$.. This is reminiscent of the concept of thermodynamic quasi-static processes, for which infinitesimal changes of external constraints take a system through a dense succession of equilibrium states [53]. Remarkably, we find that this concept generalizes to states that are even locally far from equilibrium.
@@
Conclusions. We numerically investigated the properties of a turbulent cascade arising in the GP model when a box trapped gas is periodically driven, and showed that it can be described by a universal EOS relating the turbulent state variables, the energy flux and the cascade amplitude. The form of our EOS is inconsistent with any existing theory of turbulence, posing a new theoretical challenge. Furthermore, the comparison of our classical-field simulations to experimental measurements provides valuable benchmarks for testing the validity of classical-field theories in far-from-equilibrium scenarios. Namely, we find that the GP model provides valuable insight into the experimental results of [14], even if it fails to fully capture the equation of state – indeed, the observed $na^{3}$ scaling is incompatible with the universal GP framework (see also [36]). This partial success is nevertheless striking given that there is no firm theoretical justification for its applicability in this turbulent regime. Understanding the foundations of this success, and developing refined approaches that incorporate quantum effects, remain important challenges for the field. One should also investigate the fluctuations of $N_{k}$; comparing these fluctuations between experiments and classical field simulations could shed light on the role of quantum fluctuations in turbulent quantum fluids.